OpenStudy (anonymous):

What is the solution to the equation 49^x =sqrt(7)^ (2x + 6)? a. x = 3 b. x = –3 c. x = 1 d. x = –1

OpenStudy (anonymous):

\[49^x=\sqrt{7}^{2x+6}\]\[7^{2x}=7^{x+3}\] \[2x=x+3\] \[x=3\]

OpenStudy (callisto):

\[49^x =\sqrt{7}^{2x + 6}\]\[\sqrt7^{4x} =\sqrt{7}^{2x + 6}\]4x=2x+6 2x=6 x= 6/2 =3

OpenStudy (anonymous):

why 4x?

OpenStudy (callisto):

\[49^x = 7^{2x} = \sqrt7^{2(2x)} = \sqrt7^{4x}\]

OpenStudy (anonymous):

oh i get it thanks!